# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

### v1

@@ 3 The GPR-Tree Method
A Gaussian process $f(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))$ is fully specified by its mean function $m(\mathbf{x})$ and covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$. The mean function is often assumed to be zero, $m(\mathbf{x})=0$, and the covariance function (kernel) defines the relationship between different points in the input space. A common choice for the kernel is the Matern 3/2 (M3/2) kernel, given by:
@@
Let us consider a $d$-dimensional Black-Scholes model, describer by equations (2.1). Let $N$ be the number of time steps, $\Delta t=T/N$, and $t_{n}=n\Delta t$. At any exercise date $t_{n}$, the value of the option is:
$V(t_{n},\mathbf{S}_{t_{n}})=\max\left(\Phi(S_{t_{n}}),C(t_{n},\mathbf{S}_{t_{n}})\right),$ (3.5)
@@
with $\mathbf{h}^{q}$ the $q$-th point (column vector) of the Halton sequence in $\left[0,1\right]^{d}$ and $\Phi^{-1}$ stands for the cumulative distribution function of a standarn Gaussian Variable, then applied to each component of $\mathbf{h}^{q}$. The GPR-Tree method assesses the continuation value through one step of the binomial tree proposed by Ekval [4]. For each point $\mathbf{x}^{n,p}$, we consider a set $\tilde{X}^{n,p}$ of $M=2^{d}$ possible values for $\mathbf{S}_{t_{n+1}}\mid\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$, which are computed as follows:
@@
being $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$. It is worth noticing that, as pointed out in [4], the elements of $\tilde{X}^{n,p}$ are equally likely and this simplifies the evaluation of the expected value to the computation of the arithmetic mean of the future values. Generally speaking, given $\tilde{V}_{n+1}\left(\cdot\right)$ a suitable approximation of the option value $V\left(t_{n+1},\cdot\right)$, the price function at time $t_{n}$ for $\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$ is approximated by
@@
In order to reduce the computational complexity of (3.10) when $d$ is high, see Remark 2 in Section 4. Then, GPR is employed to achive an approximation of $V\left(t_{n},\cdot\right)$ form the set $X_{n}$ of the predictors and the set $\left\{V_{n}^{Tree}\left(\mathbf{x}^{n,p}\right),p=1,\dots,P\right\}$ of the corresponding outputs. This procedure is repeated backward in time untile time $t=0$ is reached.
4 The GTU Method
Avellaneda et al. [1] propose solving the dynamic optimization problem (2.10) by dynamic programming, and here we take the same backward approach. First of all, we consider a finite number of time steps $N$, a time interval $\Delta t=T/N$ and a discrete set of times $\left\{t_{n}=n\cdot\Delta t,n=0,\dots,N\right\}$. The option price at contract inception $V\left(0,\mathbf{S}_{0}\right)$ can be computed by backward induction, starting from $t_{N}=T$. In particular, the terminal condition is given by
@@ 4 The GTU Method
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{P\in\mathcal{A}}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{S}_{t_{n+1}}\right)\mid\mathbf{S}_{t}=\mathbf{x}\right].$ (4.1)
Formally, the model parameters $P$ can vary between $t_{n}$ and $t_{n+1}$, however, following Avellaneda et al. [1], the optimal solution to the problem (4.1) can be approximated by considering a constant function with respect both time and underlying values. So, one considers the problem
@@
$D=\left[\sigma_{1}^{\min},\sigma_{1}^{\max}\right]\times\dots\times\left[\sigma_{d}^{\min},\sigma_{d}^{\max}\right]\times\left[\rho_{1,2}^{\min},\rho_{1,2}^{\max}\right]\times\dots\times\left[\rho_{d-1,d}^{\min},\rho_{d-1,d}^{\max}\right]$
$C=\left\{c_{\mathbf{x}}=\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}},\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\in D\ s.t.\ \Gamma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\succeq 0\right\}$
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{c_{\mathbf{x}}\in C}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{x}\cdot\exp\left(\left(r-\eta-\frac{\sigma^{2}}{2}\right)\Delta t+\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}}\right)^{\top}\circ\left(\Sigma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\left(\mathbf{B}_{t_{n+1}}-\mathbf{B}_{t_{n}}\right)\right)\right)\right)\right].$ (4.3)
@@
Our proposed method, the GTU algorithm, involves calculating the expected value through a single tree step, following the same approach as described in the previous Section 3. Note that this choice is particularly effective in that, once a set of parameters $c_{\mathbf{x}}$ is fixed, this calculation can be handeld as for the standard multidimensional Blasck-Scholes model:
@@
with $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$
Once we are able to compute the expected value for any $c_{\mathbf{x}}\in C$ fixed, we can tackle the optimization problem: this is a multidimensional optimisation problem with inequality constraints that are both linear (the limits for the parameters) and non-linear (the positivity condition of the $\Gamma$ matrix). This is done by means of the Sequential Quadratic Programming (SQP) algorithm, a powerful and widely used method for solving nonlinear optimization problems, particularly those with constraints. SQP methods solve a sequence of optimization subproblems, each of which approximates the original nonlinear problem by a quadratic programming problem. Drawing on the research by Biggs [2], Han [7], and Powell [10, 11], this approach enables a close simulation of Newton’s method for constrained optimization, similar to its application in unconstrained optimization. Sequential Quadratic Programming methods are considered the pinnacle of nonlinear programming techniques. For instance, Schittkowski [12] developed and evaluated a version that surpasses all other tested methods in efficiency, accuracy, and success rate across a wide array of test problems. For this reason, we have propose to use the SQP algorithm to menage problem (4.3).
@@
In order to actively implement this procedure so as to calculate the price at initial time $t=0$, one problem remains to be solved: obtaining an analytical expression of $\tilde{V}_{n}$ from observations $\tilde{V}_{n}\left(\mathbf{x}\right)$ of the same function at certain points. To do this, we use GPR regression. As described in Section 3, at each time-step $n=1,\dots,N$, we first create a mesh of points $X^{n}$ as in (3.7), but in this case the points $\mathbf{x}^{n,p}$ are define through
$\mathbf{x}_{i}^{n,p}=\mathbf{S}_{0}^{i}\exp\left(\left(r-\frac{\left(\sigma_{i}^{\text{avg}}\right)^{2}}{2}\right)t_{n}+\left(\sigma_{i}^{\text{avg}}\right)^{2}\sqrt{t_{n}}\Sigma_{i}^{\text{avg}}\Phi^{-1}\left(\mathbf{h}^{q}\right)\right),$ (4.5)
@@
Unlike neural networks, GPR requires few observations to produce good estimates of the function to be regressed, typically a few thousand. This feature combines well with the fact that, for each time-step $n$ and for each point in $X_{n}$, an optimization problem must be solved.
@@
The $\Gamma^{\text{avg}}$ matrix defined in (4.7), which is used to defined the sets $X_{n}$ for $n=1,\dots,N,$ may not be positive semidefinite. In this particular case, certain techniques can be employed to determine the closest positive semidefinite matrix to $\Gamma^{\text{avg}}$ according to the spectral norm. One of these techniques is called Projection onto the Cone of Positive Semidefinite Matrices (see e.g. Calafiore and El Ghaoui [3]). For the record, in the cases we analyzed in the numerical experiments Section (5), this situation never occurred.

### v2

@@ 3.1 GPR-Tree for option pricing
A Gaussian process $f(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))$ is fully specified by its mean function $m(\mathbf{x})$ and covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$. The mean function $m$ is often assumed to be zero, (see e.g. [Ludkovski, 2018b], [De Spiegeleer et al., 2018]), but it may also be estimated through a classical least squares regression on a set of basis functions (see e.g. [Ludkovski, 2018b], [Goudenège et al., 2021]). Furthermore, the covariance function (i.e. the kernel) defines the relationship between different points in the input space. A common choice for the kernel is the Matérn 3/2 (M3/2) kernel (see e.g. [MathWorks, 2024]), given by:
@@
Let us consider a $d$-dimensional Black-Scholes model, described by equations (2.1). Let $N$ be the number of time steps, $\Delta t=T/N$, and $t_{n}=n\Delta t$. At any exercise date $t_{n}$, the value of the option is:
$V(t_{n},\mathbf{S}_{t_{n}})=\max\left(\Psi(S_{t_{n}}),C(t_{n},\mathbf{S}_{t_{n}})\right),$ (3.5)
@@
with $\mathbf{h}^{q}$ the $q$-th point (column vector) of the Halton sequence in $\left[0,1\right]^{d}$ and $\Phi^{-1}$ stands for the cumulative distribution function of a standard Gaussian Variable, then applied to each component of $\mathbf{h}^{q}$. The GPR-Tree method assesses the continuation value through one step of the binomial tree proposed by [Ekvall, 1996]. For each point $\mathbf{x}^{n,p}$, we consider a set $\tilde{X}^{n,p}$ of $M=2^{d}$ possible values for $\mathbf{S}_{t_{n+1}}\mid\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$, which are computed as follows:
@@
being $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$. It is worth noticing that, as pointed out in [Ekvall, 1996], the elements of $\tilde{X}^{n,p}$ are equally likely and this simplifies the evaluation of the expected value to the computation of the arithmetic mean of the future values. Generally speaking, given $\tilde{V}_{n+1}\left(\cdot\right)$ a suitable approximation of the option value $V\left(t_{n+1},\cdot\right)$, the price function at time $t_{n}$ for $\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$ is approximated by
@@
In order to reduce the computational complexity of (3.10) when $d$ is high, see Remark 2 in Section 3.2. Then, GPR is employed to achieve an approximation of $V\left(t_{n},\cdot\right)$ form the set $X_{n}$ of the predictors and the set $\left\{V_{n}^{Tree}\left(\mathbf{x}^{n,p}\right),p=1,\dots,P\right\}$ of the corresponding outputs. This procedure is repeated backward in time until time $t=0$ is reached.
Remark 1.
The selection of the Matérn 3/2 kernel is based on a series of empirical tests we conducted, which also included evaluating the Matérn 5/2 kernel and the Squared Exponential kernel. From these numerical experiments, we observed that, while there are no significant differences in the numerical results among the tested kernels, the best overall performance was achieved with the Matérn 3/2 kernel. Generally, the proposed kernel is particularly well-suited for modeling financial option prices using Gaussian Process Regression due to its ability to balance flexibility and smoothness. Unlike smoother kernels such as the Matérn 5/2 or Squared Exponential kernels, the Matérn 3/2 kernel captures local variations and abrupt changes more effectively, which are often present in the relationship between option prices and underlying asset values. Finally, we note that similar findings about these kernels have been reported by [Ludkovski, 2018b] and [Goudenège et al., 2020].
Remark 2.
In Gaussian-Process regression, the characteristic length-scale $\sigma_{l}$, the signal variance $\sigma_{f}^{2}$, and the observation-noise variance $\sigma_{n}^{2}$ are termed hyperparameters. Following the convention adopted in Matlab’s fitrgp routine, we estimate these quantities by maximizing the log-marginal likelihood of the training data, an approach equivalent to Type-II maximum-likelihood (empirical Bayes) inference (for more details see [Williams and Rasmussen, 2006] and [MathWorks Inc., 2024]).
3.2 GTU: GPR-Tree for the UVM
[Avellaneda et al., 1995] propose solving the dynamic optimization problem (2.10) by dynamic programming, and here we take the same backward approach. First of all, we consider a finite number of time steps $N$, a time interval $\Delta t=T/N$ and a discrete set of times $\left\{t_{n}=n\cdot\Delta t,n=0,\dots,N\right\}$. The option price at contract inception $V\left(0,\mathbf{S}_{0}\right)$ can be computed by backward induction, starting from $t_{N}=T$. In particular, the terminal condition is given by
@@ 3.2 GTU: GPR-Tree for the UVM
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{\boldsymbol{\theta}\in\Theta}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{S}_{t_{n+1}}\right)\mid\mathbf{S}_{t}=\mathbf{x}\right].$ (3.11)
Formally, the model parameters $P$ can vary between $t_{n}$ and $t_{n+1}$, however, following [Avellaneda et al., 1995], the optimal solution to the problem (3.11) can be approximated by considering a constant function with respect both time and underlying values. So, one considers the problem
@@
$D=\left[\sigma_{1}^{\min},\sigma_{1}^{\max}\right]\times\dots\times\left[\sigma_{d}^{\min},\sigma_{d}^{\max}\right]\times\left[\rho_{1,2}^{\min},\rho_{1,2}^{\max}\right]\times\dots\times\left[\rho_{d-1,d}^{\min},\rho_{d-1,d}^{\max}\right],$
$\Theta_{\mathbf{x}}=\left\{\boldsymbol{\theta}_{\mathbf{x}}=\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}},\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\in D\ s.t.\ \Gamma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\succeq 0\right\},$
where $\Gamma\left(\cdot\right)\succeq 0$ means that the $\Gamma\left(\cdot\right)$ must be a positive semidefinite matrix. Moreover, we set
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{\boldsymbol{\theta}_{\mathbf{x}}\in\Theta_{\mathbf{x}}}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{x}\cdot\exp\left(\left(r-\eta-\frac{\sigma^{2}}{2}\right)\Delta t\right.\right.\right.\\ \left.\left.\left.+\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}}\right)^{\top}\circ\left(\Sigma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\left(\mathbf{B}_{t_{n+1}}-\mathbf{B}_{t_{n}}\right)\right)\right)\right)\right].$ (3.13)
@@
Our proposed method, the GTU algorithm, involves calculating the expected value through a single tree step, following the same approach as described in the previous Section 3.1. Note that this choice is particularly effective in that, once a set of parameters $\boldsymbol{\theta}_{\mathbf{x}}$ is fixed, this calculation can be handled as for the standard multidimensional Black-Scholes model:
@@
with $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$ We emphasize that, in formula (3.14), the function $V_{n+1}$ is evaluated at $2^{d}$ points, aiming to discretize the possible values of the random variable $S_{t_{n+1}}\mid S_{t_{n}}$ by approximating the Gaussian vector $\mathbf{B}_{t_{n+1}}-\mathbf{B}_{t_{n}}$ with a discrete random vector, uniformly distributed in $\{-1,+1\}^{d}$, following the approach proposed by [Ekvall, 1996].
Once we are able to compute the expected value for any $\boldsymbol{\theta}_{\mathbf{x}}\in\Theta_{\mathbf{x}}$ fixed, we can tackle the optimization problem: this is a multidimensional optimisation problem with inequality constraints that are both linear (the limits for the parameters) and non-linear (the positivity condition of the $\Gamma$ matrix). This is done by means of the Sequential Quadratic Programming (SQP) algorithm, a powerful and widely used method for solving nonlinear optimization problems, particularly those with constraints. SQP methods solve a sequence of optimization subproblems, each of which approximates the original nonlinear problem by a quadratic programming problem. Drawing on the research by [Biggs, 1975], [Han, 1977], and [Powell, 1978, Powell, 2006], this approach enables a close simulation of Newton’s method for constrained optimization, similar to its application in unconstrained optimization. Sequential Quadratic Programming methods are considered the pinnacle of nonlinear programming techniques. For instance, [Schittkowski, 1986] developed and evaluated a version that surpasses all other tested methods in efficiency, accuracy, and success rate across a wide array of test problems. For this reason, we have propose to use the SQP algorithm to manage problem (3.13).
@@
In order to actively implement this procedure so as to calculate the price at initial time $t=0$, one problem remains to be solved: being able to estimate $\tilde{V}_{n}$ at any significant point by exploiting the observations $\tilde{V}_{n}\left(\mathbf{x}\right)$ of the same function at certain points. This is important because the calculation of $V_{n-1}$ relies on the calculation of the function $V_{n}$ at appropriate points. To this aim, we use GPR regression. As described in Section 3.1, at each time-step $n=1,\dots,N$, we first create a mesh of points $X^{n}$ as in (3.7), but in this case the points $\mathbf{x}^{n,p}$ are define through
$\mathbf{x}_{i}^{n,p}=\mathbf{S}_{0}^{i}\exp\left(\left(r-\frac{\left(\sigma_{i}^{\text{avg}}\right)^{2}}{2}\right)t_{n}+\sigma_{i}^{\text{avg}}\sqrt{t_{n}}\Sigma_{i}^{\text{avg}}\Phi^{-1}\left(\mathbf{h}^{q}\right)\right),$ (3.15)
@@
Unlike NNs, GPR requires few observations to produce good estimates of the function to be regressed, typically a few thousand (to this aim, see e.g. [Kamath et al., 2018]) and [Mallick et al., 2021] ). This feature combines well with the fact that, for each time-step $n$ and for each point in $X_{n}$, an optimization problem must be solved.
@@
The $\Gamma^{\text{avg}}$ matrix defined in (3.17), which is used to defined the sets $X_{n}$ for $n=1,\dots,N,$ may not be positive semidefinite. In this particular case, certain techniques can be employed to determine the closest positive semidefinite matrix to $\Gamma^{\text{avg}}$ according to the spectral norm. One of these techniques is called Projection onto the Cone of Positive Semidefinite Matrices (see e.g. [Calafiore and El Ghaoui, 2014]). For the record, in the cases we analyzed in the numerical experiments Section (5), this situation never occurred.
4 The NNU method
4.1 Neural Networks
Neural networks (NNs) are a class of parametric functions widely used in ML due to their remarkable ability to approximate complex mappings. From a mathematical standpoint, a neural network is a function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ constructed as a composition of affine maps and nonlinear activation functions. More precisely, for a given depth $L\in\mathbb{N}$, and layer widths $(d_{0},d_{1},\dots,d_{L})$, a (feed-forward) NN is defined by
$F(x)=W_{L}\circ\phi\circ W_{L-1}\circ\cdots\circ\phi\circ W_{1}(x),$
where each $W_{\ell}(x)=A_{\ell}x+b_{\ell}$ is an affine transformation with weight matrix $A_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and bias vector $b_{\ell}\in\mathbb{R}^{d_{\ell}}$, and $\phi:\mathbb{R}\to\mathbb{R}$ is a fixed nonlinear activation function applied componentwise.
This architectural structure allows NNs to form highly flexible families of functions. The remarkable approximation capacity of NNs is rooted in their expressive power. Under mild assumptions on the activation function $\phi$, it can be shown that NNs with a single hidden layer are capable of approximating any continuous function on compact subsets of $\mathbb{R}^{d}$ arbitrarily well. Specifically, if $\phi$ is continuous, bounded, and non-constant, then for every continuous function $f:K\to\mathbb{R}$ defined on a compact set $K\subset\mathbb{R}^{d}$, and for every $\varepsilon>0$, there exists a NN $F$ with one hidden layer and a finite number of units such that
$\sup_{x\in K}|f(x)-F(x)|<\varepsilon.$
This result, known as the Universal Approximation Theorem and developed by [Hornik, 1991], provides the theoretical foundation for the use of NNs in function approximation tasks. In particular, it justifies their application in high-dimensional problems arising in quantitative finance, where NNs can approximate hedging strategies, pricing functionals, and conditional expectations under complex market dynamics.
This result implies that NNs can approximate any continuous function to arbitrary precision, provided the network is sufficiently wide. In practice, this universality – along with advances in optimization and hardware – has led to the successful application of NNs in high-dimensional function approximation tasks, including those arising in finance, such as hedging, pricing, and risk management. We refer the interested reader to [Buehler et al., 2019], where some recent applications of NNs in finance are discussed.
4.2 NNU: Neural Networks for UVM
In [Buehler et al., 2019], the authors introduce the Deep Hedging framework, where feed-forward NNs are employed to solve a stochastic control problem: determining the optimal hedging strategy that minimizes a suitable penalty function evaluating the effectiveness of hedging. The strategy is learned by feeding the NN a large number of simulated market states and training the NN to output the quantity of each instrument to include in a hedging portfolio associated with a short position on a given derivative.
In our work, we adopt and reformulate these principles to address a different problem: given a state of the market, we aim to determine the volatilities with which to propagate risky assets so as to maximize (or minimize) the expected value of a terminal payoff. The underlying optimization problem is structurally similar, but the control variables shift from hedging positions to volatility parameters, and the performance objective changes accordingly to reflect our targeted financial quantity. It is worth observing that, unlike Deep Hedging – which optimizes trading strategies under frictions using semi-recurrent networks and convex risk measures – the proposed NNU approach employs a compact feedforward NN that directly maps each market state, along with the time of market data observation, to worst-case volatilities and correlations. Although both methods leverage NNs, they differ fundamentally in application domain and network design-highlighting the originality of NNU in addressing control problems under parameter uncertainty rather than dynamic hedging under trading constraints.
4.3 NNU under fixed correlation
Let us now discuss in detail how the NNU algorithm works when correlation is certain, and thus matrix $\Sigma$ is fixed. First of all, as done in Section 3.2, we consider a finite number of time steps $N$, a time interval $\Delta t=T/N$ and a discrete set of times $\left\{t_{n}=n\cdot\Delta t,n=0,\dots,N\right\}$. As a first step, $M$ random increments
$\left\{\Delta\mathbf{B}_{n,m}=\left(B^{1}_{t_{n},m}-B^{1}_{t_{n-1},m},\dots,B^{d}_{t_{n},m}-B^{d}_{t_{n-1},m}\right)^{\top},n=1,\dots,N,m=1,\dots,M\right\}$
of the $d$-dimensional uncorrelated Wiener processes $\mathbf{B}$, are sampled for the $N$ time increments. Then, by using those increments, the network generates $M$ trajectories for the prices of the underlying asset, denoted by
$\left\{\mathbf{S}_{t_{n},m}=\left(S^{1}_{t_{n},m},\dots,S^{d}_{t_{n},m}\right)^{\top},n=0,\dots,N,m=1,\dots,M\right\},$
with $\mathbf{S}_{0,m}=\mathbf{S}_{0}$ for all trajectories. Specifically, these values are generated through the equation
$\mathbf{S}_{t_{n+1},m}=\mathbf{S}_{t_{n},m}\circ\exp\left(\left(r-\boldsymbol{\eta}-\frac{1}{2}F(t_{n},\mathbf{S}_{t_{n},m})^{2}\right)\Delta t+F(t_{n},\mathbf{S}_{t_{n},m})\circ\left(\Sigma\,\Delta\mathbf{B}_{n+1,m}\right)\right),$ (4.1)
[…]
$AP=\frac{e^{-rT}}{M}\sum_{m=1}^{M}\Psi\left(\mathbf{S}_{t_{N},m}\right).$
The network parameters are optimized during training with the objective of maximizing $AP$, which, upon completion of the training process, is taken as an estimator of the contract’s initial price.
The network architecture consists of batch normalization, dense layers with ReLU activation, dropout regularization, and a final sigmoid layer scaled to the interval $[\sigma_{i}^{\min},\sigma^{\max}_{i}]$ for each volatility. Moreover, since the algorithm relies on Monte Carlo simulations to approximate the expected value of the payoff, it is also possible to estimate confidence intervals for the price.
At the end of training, the learned model defines a mapping from market states to optimal volatilities, effectively controlling the stochastic evolution of asset prices to achieve a target objective. This represents a novel application of NNs to model uncertainty in financial systems through learned volatility control.
4.4 NNU under correlation uncertainty
The NNU algorithm can also be applied to handle uncertainty in the correlation structure among the underlying assets. In principle, the algorithm should be adapted so that, in addition to volatilities $\boldsymbol{\sigma}$, it also returns correlations $\left\{\rho_{i,j},1\leq i<j\leq d\right\}$ that define the matrix $\Gamma$ as in (2.4), from which it derives – by Cholesky factorization – the matrix $\Sigma$ employed in (4.1).
In the case of dimension $d=2$, the implementation of such an update is relatively straightforward; the addition of an output to the network suffices. However, in higher dimensions, the process becomes more intricate because, when the network outputs a vector of tentative correlations – despite the fact that they are bounded values – the corresponding correlation matrix may fail to be positive semidefinite, resulting in the breakdown of the entire procedure. To guarantee well-posed simulations we employ an external-penalty method (see e.g. [Nocedal and Wright, 2006] and [Fletcher, 2013]) for the constrained optimisation problem
$\max_{\mathbf{\theta}\in\mathbf{\Theta}}\;\frac{e^{-rT}}{M}\sum_{m=1}^{M}\Psi\left(\mathbf{S}_{t_{N},m}\right)\quad\text{subject to }\Gamma\left(t_{n},\mathbf{S}_{t_{n},m}\right)\succeq 0\;\forall n\text{ and }m.$
Specifically, given the raw network output $\widehat{\Gamma}\left(t,\mathbf{S}_{t_{n},m}\right)$, we actually employ as correlation matrix
$\Gamma\left(t,\mathbf{S}_{t_{n},m}\right)\;=\;\begin{cases}\widehat{\Gamma}\left(t,\mathbf{S}_{t_{n},m}\right),&\text{if }\widehat{\Gamma}\left(t,\mathbf{S}_{t_{n},m}\right)\succeq 0,\\[4.0pt] \Gamma_{0},&\text{otherwise},\end{cases}$ (4.2)
where $\Gamma_{0}$ is an admissible matrix with respect to the constraints on the correlations (in our numerical test we set $\Gamma_{0}$ as the identity matrix $I_{d}$). Thanks to equation (4.2), the forward pass (4.1) always uses a valid correlation matrix and the Cholesky factor exists. The map $\widehat{\Gamma}\mapsto\Gamma$ acts as a crude projection onto the interior of the cone of symmetric positive semidefinite matrices.
Exiting the set of positive semidefinite matrices is passively penalized (we force the algorithm to use the $\Gamma_{0}$ matrix that is not necessarily optimal) but also actively through the introduction of a penalty term that discourages the network from generating ineligible matrices. Specifically, let us define the Frobenius-distance penalty
[…]
with $\|\cdot\|_{Fr}$ the Frobenius norm. For a fixed $\lambda>0$ we optimise the penalised objective
$\frac{e^{-rT}}{M}\sum_{m=1}^{M}\Psi\left(\mathbf{S}_{t_{N},m}\right)+\lambda\sum_{m=1}^{M}\sum_{n=0}^{N-1}\pi\left(\Gamma\left(t_{n},\mathbf{S}_{t_{n},m}\right)\right).$
This penalty mechanism is effective for several intertwined mathematical reasons. First, by construction, it ensures that the simulation of the paths is always well defined and respectful of the constraints. Whenever the network proposes a matrix that is not positive semidefinite, we fall back to the admissible matrix $\Gamma_{0}$, a valid correlation matrix for the optimization problem. This guarantees that the Cholesky factor exists and avoids any runtime breakdown in the simulation. In addition, the UVM model’s constraints on correlations are met.
From the perspective of optimisation theory, the method can be viewed as an external penalty formulation of the constrained optimisation problem. As shown in classical references such as [Nocedal and Wright, 2006] and [Fletcher, 2013], this type of approach ensures that, under mild assumptions (e.g., continuity and boundedness of the objective function), minimisers of the penalised objective converge to minimisers of the original constrained problem as the penalty weight $\lambda$ tends to infinity.
An important advantage of this approach lies in the stability of the training process. Since the forward simulation step never leaves the differentiable domain of the Cholesky factor, the gradients computed during backpropagation remain finite, avoiding instabilities. Furthermore, the penalty function $\pi(\cdot)$, defined via a Frobenius distance to the identity, is Lipschitz continuous and differentiable almost everywhere, which makes it well suited for stochastic gradient methods.
Empirically, the mechanism has the beneficial effect of steering the network towards feasible outputs very early in training. From numerical experiments, we observed that, after a few epochs, the network produces positive semidefinite correlation matrices almost everywhere, at which point the penalty becomes inactive and the optimisation proceeds as if constrained directly on the target objective. Hence the proposed penalty mechanism provides a computationally cheap safeguard that keeps the simulation in the feasible set, delivers stable learning dynamics, and – for $\lambda$ large enough – drives the network towards admissible correlation structures without altering the original objective.